# A new economic and financial theory of money

### v6

@@ I Introduction
Microeconomic theory and macroeconomic theory were developed independent of each other, and a weak attempt was made to understand the relationship between them, unlike the strong attempt made in the physical systems example shown in Fig. 1. Inspired by the television series Stargate and Mr. Robot, we will proceed by identifying macroeconomics as the uber theory. We propose that, not only the sovereign, but all sub-economies and entities should be motivated to maximize the value of currencies that are relevant to them (those currencies in which they economically transact). With the advent of electronic currencies, there now can be currencies for each sub-economy whether that be regional, industry-based or both. The relationship of these currencies to each other defines a graph of the economy, and the monetary and fiscal policy of the individual currencies give economic control knobs matched to the structure (mathematically, the topology) of the economy. Instead of a constrained maximization of an entity’s DCF or NPV (that is, short-term profit) with the constraint of $\text{NPV}=\text{DCF}-I>0$, the entity should be motivated to maximize the long-term virtuous economic activity of the whole economy $m_{e}S_{0}R_{0}$ with the constraint of not inflating the value of the currency (that is $m_{e}S_{0}R_{0}-I>0$). This economic theory will be discussed in more detail in Sec. III, Sec. VII, Sec. VIII, Sec. X and Sec. IX with a detailed example of how it works given in Sec. XII, Sec. XIII and App. A.
@@
We return now to the concept of economic conservatism embodied in the seminal work of Milton Friedman and summarized in the book “Capitalism & Freedom” (Friedman, 2020). Like Friedman advocated, Government Inc. should be limited to matters of the common good like defense, education, social security, public infrastructure and resources, the arts, fundamental research and healthcare. There should be freedom, not only of sub-economies, but of Government Inc., labor, intellectual property and law. This will be discussed in more detail in Sec. XV. Friedman was right that freedom is essential. Freedom is how the leaders of the sub-economies are ultimately held accountable. Free people will move from a sub-economy with a poor leader to a sub-economy with a good leader. The problem comes with capitalism being based on the short-term maximization of profits. The fact that it is left up to the private sector (the second ingredient of capitalism) is common to what we are proposing. It would be best if Government Inc. is private with government executives and congressional representatives serving on the board. The change that is being proposed is to move to “Communism & Freedom”, where we are talking about a private sector communism where the virtuous economic activity (that is the common good) is maximized. It is very similar to existing economic cooperatives, but with a currency for the economic cooperative.
@@ II Catastrophic consequences of the invalid economic localization
This section will start by examining the local diffusive approximation resulting in Eq. 4 and Eq. 1, what it really means, how it becomes embedded in our economy, and why it is not a valid approximation for many cases, especially for infrastructure and other long-term capital assets. The primary economic consequences of using this approximation will then be explored, followed by the secondary economic consequences, and finally the effects on the fabric of society.
@@
where $H(J)$ is the Hamiltonian or energy as a function of $J$ with units of $J/t$ (if $J$ is physical action, then $H(J)$ has units of energy), $\omega(J)=dH/dJ$, and $T$ is a constant temperature with units of $J/t$. It will be shown in Sec. VII that $\kappa=\nu\,\sigma_{c}^{2}\sim T^{3/2}$, $\nu\sim\sqrt{T}$, and $\sigma_{c}^{2}\sim T$. So what is identified by the temperature $T$, or the random kinetic energy, of a heat bath or an external physical system, is proportional to the 2/3 root of the diffusion coefficient of a random walk or diffusive process. For financial dynamics the diffusion coefficient, $\kappa=\left<\Delta P\Delta P/\Delta t\right>$, is called the volatility (which is the metric of risk), where $P$ is the price of the commodity. Equation 6, when $P$ is substituted for $J$, is the Black-Scholes equation of quantitative finance.
This assumption is made because it dramatically simplifies the general problem from the difficult to solve Hamilton-Jacobi-Bellman equation (Eq. 37), to an easily solvable partial differential equation (Eq. 6) with a closed form, time independent (stationary), equilibrium solution given by Eq. 7. It is simple as foretold by Murray Gell-Mann.
@@ III Theory and valuation of electronic currencies
is the ensemble and long time average revenue of the primary sub-economy expressed as a functional integral over the distribution functional $f[R(t)]$ and time. Equation 8 can be summed or integrated to give
@@
is the effective economic multiplier. Equation 16 is what we call the fundamental equation of electronic currency valuation. The relationship to the value of the electronic currency can be seen by writing
@@
which is Eq. 3 in Sec. I with more precise definitions.
Let us now propose that the electronic currency management firm should maximize the value of the electronic currency, $P_{\text{ec}}$, as given by Eq. 20. This is in the interest of its stakeholders – the holders of the electronic currency. Equation 16 is made up of four parts.
@@
where $\Delta R_{0}$ is the increase in revenue coming from the investment financed by issuing $\Delta I$ in new electronic currency, and $\Delta M$ is the increase in monetary demand net of the investment, $\Delta M\equiv m_{e}S_{0}\Delta R_{0}-\Delta I$. Note that the costs do not appear explicitly in this investment criteria (a small fraction of the costs appear in the required investment), unlike the $\text{NPV}>0$ constraint where all the costs appear explicitly as $E$ in Eq. 1. The revenue will be adjusted so that it includes all non-monetary value and/or mitigated costs to society. This will be implemented as a tax on undesirable activities or a subsidy of desirable activities. If the non-monetary value and/or mitigated costs are felt by members outside of the sub-economy, there should be a taxation of the members outside of the sub-economy for undertaking undesirable activities and investment of the proceeds from that taxation into the sub-economy, that is a subsidy of the desirable activities, by the government. Note that if the tax is sufficiently large, the government or the currency management firm will never have to collect the tax. The undesirable activities will not be undertaken.
@@
Now, identify the two boundaries given by the conditions $\Delta M^{\prime}(Q_{\text{max}})=0$ and $R^{\prime}(Q_{\text{min}})=0$. Using Eq. 31, find that $\lambda_{\text{max}}=\infty$ and $\lambda_{\text{min}}=0$, and using Eq. 28, find that $\mu_{\text{max}}=\Delta M(Q_{\text{max}})$ and $\mu_{\text{min}}=\Delta M(Q_{\text{min}})$. Finally, summarize the solution for the different values of $\mu$. It is not possible for $\mu$ to be greater than $\mu_{\text{max}}$. If $\mu=\mu_{\text{max}}$
@@
The electronic currency management firm will want to maximize $\mu$ in order to maximize the value of the currency, so that $\mu=\mu_{\text{max}}$. Therefore, the solution is given by Eq. 32. For a well managed currency, $m_{e}S_{0}/T_{I}\gg 1$. For instance, the example given in App. A has $T_{I}=1\,\text{yr}$, $S_{0}=2\,\text{yr}$ and $m_{e}=3.5$, so that $m_{e}S_{0}/T_{I}=7$ and $R^{\prime}=(1/8)E^{\prime}$. This is close to being a revenue maximizing firm with $R^{\prime}=0$ and $\mu=\mu_{\text{min}}$, which is not much less than $\mu_{\text{max}}$ since
@@
for the example given in App. A. In contrast, if the local approximation is made, the value of the currency is still maximized by using Eq. 32, but the non-local business network and time effects are neglected. The effective economic multiplier $m_{e}$ will approach 1, and the savings multiplier $S_{0}$ will approach 0. For Target Energy Inc., the typical firm operated with the local assumption discussed in App. A, $m_{e}=2$ and $S_{0}=1/20\,\text{yr}$. This gives $m_{e}S_{0}/T_{I}=1/10$, so that $R^{\prime}=(10/11)E^{\prime}\approx E^{\prime}$ – a profit, not revenue, maximizing firm.
The functional integration over the distribution given by the functional $f[R(t)]$ in Eq. 9 is how the risk model enters this theory. This is what could be approximated by the Generative Pretrained Transformer (GPT) (Li et al., 2022), Generative Adversarial Network (GAN) (Goodfellow et al., 2016), or Heisenberg Scattering Transformation (HST) (Glinsky and Maupin, 2023; Glinsky, 2024a, 2023a, 2023b), all forms of Artificial Intelligence (AI) (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016). In the case of the local approximation, it is approximated by solving the diffusive Black-Scholes equation given in Eq. 6. There are also more simple approximations that can be made that respect the low risk and long-term nature of many financial situations.
Starting with the equation for $R_{0}$, Eq. 9, and substituting the solutions to the Black-Scholes equation, Eq. 6, into it, one can derive the expression for NPV given in Eq. 1. It will be shown in Sec. VII and Sec. IX how a constrained optimization of the expression for $R_{0}$ given in Eq. 9 leads to the optimization of NPV given in Eq. 1, when a dominate diffusive risk model is superimposed on the conservative financial dynamics of a sub-economy.
The dramatic difference between using the true distribution functional $f[R(t)]$ of Eq. 9 and the distribution function $f(J,t)$ of Eq. 6 is demonstrated in Fig. 3. In this figure the HST and the methods of AI are used to conditionally generate the distribution based on the historical behavior of the oil price (as will be explained in detail in Sec. VII). Note the discounting of the value of the next business cycle by the local approximation, so that no inventory would be amassed to take advantage of the next market upswing. Also note how poor the simulations using the local approximations are – that is how bad the approximation is.
Figure 3: Examples of conditional generation of price profile $R(t)$. (a) using the distribution functional $f[R(t)]$ of Eq. 9 via the HST, (b) using the distribution function $f(R,t)$ of Eq. 6. The thick black line is the actual signal. The thin black lines are the conditional realizations.
@@
In Sec. VII, it will be shown how to solve Eq. 37 using the HST and the methods of AI based on either simulations of the system or observations of the system, driven by an external force $F_{\text{ext}}(p,q)$. The solution will be the fiscal (investment and operational) and monetary policy $\pi^{*}(q)$ that will locally maximize $V(q,t)$, but needs to be stabilized and cooled since local maximums are unstable equilibriums. The reward that is optimized $R(q)$ can be modified to take into account other benefits or costs to society, such as sustainability and beauty, by applying a conservative economic control force $\Delta p/\Delta t=F_{c}(q)$. Furthermore, the application of the economic force (arbitrage trading) to stabilize the system equilibrium $\pi^{*}(q)$ and minimize the fluctuations about the system equilibrium can be done in a similarly direct feedback manner by applying a force $F_{sf}(P)$ given in Eq. 99. It also can be done by applying a ponderomotive and diffusive force $F_{sp}(t)$, given in Eq. 100, that is not dependant on knowing the state of the system. More details of these control forces can be found in Sec. VII.
If instead of the true model of the system, the diffusive model of Eq. 6 is used and applied to solve Eq. 37, it can be shown that the equilibrium optimum policy $\pi^{*}(q)$ is to optimize the NPV of Eq. 1 and to invest with terms of debt repayment. This is equivalent to the profit maximization of the previous constrained optimization, instead of a revenue maximization. It should be noted that the local NPV maximization leads to a significant reduction in the long-term economic activity and even long-term profits. This greedy, local optimization when the system does not have exponential risk (that is diffusive dynamics), leads to very poor strategy (that is, policy). It would be like playing chess with no regard for the long-term effects of the next move. One would never sacrifice a piece to improve the long-term prospects of winning. One would only consider what pieces could be captured or lost with the next move or two. It also would be like playing PacMan with a strategy to swallow the next cherry as quickly as possible with no concern to whether that path leads to being consumed by a ghost in the near future.
@@ V Management of sub-economy
The criteria on which investments should be made is whether the demand for the electronic currency that the investment will create is greater than the investment, as shown in Eq. 22. The electronic currency management firm can then electronic currency leverage the funds that it has in reserves by issuing new electronic currency to make the investment. This ensures a future growth in the value of the currency, and that the sub-economy will not be paying for the investment through inflation, effectively an inflation tax. This is in contrast to issuing debt, which exploits the sub-economy and leads to sub-optimal levels of investment and economic performance of the sub-economy.
@@ VI Sub-economy as a nonlinear system
In this section, we will approach the understanding of a financial system as a nonlinear system, and the financial and monetary policies as a problem in nonlinear systems control. We start this discussion by referring to Fig. 5, which is a representative contour plot for a physical system (Kuzmin et al., 2004). It also can be looked at as a topography map. This map has two basins with basin centers indicated by the o-points, and one saddle point (mountain pass) between the basins indicated by the x-point. There are strong exploitative thermal forces that will take a sub-economy to the o-point once in the respective basin. They are also what are called stable equilibriums. The discipline of nonlinear dynamics refers to these o-points or stable equilibriums as attractors. This is in contrast to the x-point, which is the point that the sub-economy descends to from the mountain, but it is not a stable equilibrium so that as it is approached the sub-economy will fall into one of the exploitative basins. The discipline of nonlinear dynamics refers to these x-points or unstable/metastable equilibriums as semi-attractors (that is half attractor and half repeller). Semi-attractors first attract the trajectory to them, but once reached repel the trajectory away from them. These unstable equilibriums are the desirable states from a social aesthetic perspective as shown in Fig. 6. They need an active control system, though, in order to stabilize them. This is like stabilizing an inverted pendulum with a vibrating saw as shown in Fig. 7 (Kapitza, 1951a, b; Landau and Lifshitz, 1976). Stabilization is the construction of a small alpine valley at the mountain pass where it is easiest to do. The electronic currency management firm needs to vibrate the market with its arbitrage trading to stabilize the system. The technical problem is in identifying the natural frequencies of the system. For instance, when one flies a plane, one must take into account that it takes several seconds for the plane to respond. If one tries to make corrections faster than this, one will over-correct and cause the plane to go out of control. The issue with financial systems is that they have many natural frequencies, in fact an infinite number. The electronic currency management firm needs to use innovative genAI technology that will control the financial system at all the natural frequencies. Details of how this is done are given in the US Patent Application, “Systems and Methods for Controlling and Simulating Complex Systems” (Glinsky, 2023c) and discussed in Sec. VII. Current control systems only have a single time scale, and often have catastrophic phase lags built in that destabilize the system like proof-of-stake systems (e.g., New Ethereum) and the buying and selling of bonds by a central bank (e.g., US T-bills and British GILTs).
@@
Figure 7: Stabilization of the pendulum about the unstable equilibrium (saddle point) by use of nonlinear control. The full video can be found at this Link.
Note that the electronic currency management firm will want to buy the electronic currency, using its cash reserves, when the price is less than the equilibrium price to increase the demand and therefore the price, and sell the electronic currency when the price is more than the equilibrium price to increase the supply and therefore decrease the price. That is, it will buy low and sell high – a money making financial heat engine. Mother nature rewards doing what she wants. The essential tricks are sensing what the equilibrium is and doing it on multiple time scales. It is all about topological discovery of the financial system response and equilibrium, then topological manipulation to control and stabilize the financial system (Glinsky, 2024b, c). This is not the case today.

### v7

@@ I Introduction
Microeconomic theory and macroeconomic theory were developed independent of each other, and a weak attempt was made to understand the relationship between them, unlike the strong attempt made in the physical systems example shown in Fig. 1. Inspired by the television series Stargate and Mr. Robot, we will proceed by identifying macroeconomics as the uber theory. We propose that, not only the sovereign, but all sub-economies and entities should be motivated to maximize the value of currencies that are relevant to them (those currencies in which they economically transact). With the advent of electronic currencies, there now can be currencies for each sub-economy whether that be regional, industry-based or both. The relationship of these currencies to each other defines a graph of the economy, and the monetary and fiscal policy of the individual currencies give economic control knobs matched to the structure (mathematically, the topology) of the economy. Instead of a constrained maximization of an entity’s DCF or NPV (that is, short-term profit) with the constraint of $\text{NPV}=\text{DCF}-I>0$, the entity should be motivated to maximize the long-term virtuous economic activity of the whole economy $m_{e}S_{0}R_{0}$ with the constraint of not inflating the value of the currency (that is $m_{e}S_{0}R_{0}-I>0$). This economic theory will be discussed in more detail in Sec. III, Sec. VII, Sec. VIII, Sec. IX, and Sec. X with an example of how it works given in Sec. XII (details of which appear in App. A). The practical new Ubuntu business model resulting from this economic theory, replacing the conventional debt based business model, is described in Sec. XIII.
@@
We return now to the concept of economic conservatism embodied in the seminal work of Milton Friedman and summarized in the book “Capitalism & Freedom” (Friedman, 1962). Like Friedman advocated, Government Inc. should be limited to matters of the common good like defense, education, social security, public infrastructure and resources, the arts, fundamental research and healthcare. There should be freedom, not only of sub-economies, but of Government Inc., labor, intellectual property and law. This will be discussed in more detail in Sec. XV. Friedman was right that freedom is essential. Freedom is how the leaders of the sub-economies are ultimately held accountable. Free people will move from a sub-economy with a poor leader to a sub-economy with a good leader. The problem comes with capitalism being based on the short-term maximization of profits. The fact that it is left up to the private sector (the second ingredient of capitalism) is common to what we are proposing. It would be best if Government Inc. is private with government executives and congressional representatives serving on the board. The change that is being proposed is to move to “Communism & Freedom”, where we are talking about a private sector communism where the virtuous economic activity (that is the common good) is maximized. It is very similar to existing economic cooperatives, but with a currency for the economic cooperative.
@@ II Catastrophic consequences of the invalid economic localization
This section will start by examining the local diffusive approximation resulting in Eq. (4) and Eq. (1), what it really means, how it becomes embedded in our economy, and why it is not a valid approximation for many cases, especially for infrastructure and other long-term capital assets. The primary economic consequences of using this approximation will then be explored, followed by the secondary economic consequences, and finally the effects on the fabric of society.
@@
where $H(J)$ is the Hamiltonian or energy as a function of $J$ with units of $J/t$ (if $J$ is physical action, then $H(J)$ has units of energy), $\omega(J)=dH/dJ$, and $T$ is a constant temperature with units of $J/t$. It will be shown in Sec. VII that $\kappa=\nu\,\sigma_{c}^{2}\sim T^{3/2}$, $\nu\sim\sqrt{T}$, and $\sigma_{c}^{2}\sim T$. So what is identified by the temperature $T$, or the random kinetic energy, of a heat bath or an external physical system, is proportional to the 2/3 root of the diffusion coefficient of a random walk or diffusive process. For financial dynamics the diffusion coefficient, $\kappa=\left<\Delta P\Delta P/\Delta t\right>$, is called the volatility (which is the metric of risk), where $P$ is the price of the commodity. Equation (6), when $P$ is substituted for $J$, is the Black-Scholes equation of quantitative finance.
This assumption is made because it dramatically simplifies the general problem from the difficult to solve Hamilton-Jacobi-Bellman equation (Eq. (37)), to an easily solvable partial differential equation (Eq. (6)) with a closed form, time independent (stationary), equilibrium solution given by Eq. (7). It is simple as foretold by Murray Gell-Mann.
@@ III Theory and valuation of electronic currencies
is the ensemble and long time average revenue of the primary sub-economy expressed as a functional integral over the distribution functional $f[R(t)]$ and time. Equation (8) can be summed or integrated to give
@@
is the effective economic multiplier. Equation (16) is what we call the fundamental equation of electronic currency valuation. The relationship to the value of the electronic currency can be seen by writing
@@
which is Eq. (3) in Sec. I with more precise definitions.
Let us now propose that the electronic currency management firm should maximize the value of the electronic currency, $P_{\text{ec}}$, as given by Eq. (20). This is in the interest of its stakeholders – the holders of the electronic currency. Equation (16) is made up of four parts.
@@
where $\Delta R_{0}$ is the increase in revenue coming from the investment financed by issuing $\Delta I$ in new electronic currency, and $\Delta M$ is the increase in monetary demand net of the investment, $\Delta M\equiv m_{e}S_{0}\Delta R_{0}-\Delta I$. Note that the costs do not appear explicitly in this investment criteria (a small fraction of the costs appear in the required investment), unlike the $\text{NPV}>0$ constraint where all the costs appear explicitly as $E$ in Eq. (1). The revenue will be adjusted so that it includes all non-monetary value and/or mitigated costs to society. This will be implemented as a tax on undesirable activities or a subsidy of desirable activities. If the non-monetary value and/or mitigated costs are felt by members outside of the sub-economy, there should be a taxation of the members outside of the sub-economy for undertaking undesirable activities and investment of the proceeds from that taxation into the sub-economy, that is a subsidy of the desirable activities, by the government. Note that if the tax is sufficiently large, the government or the currency management firm will never have to collect the tax. The undesirable activities will not be undertaken.
@@
Now, identify the two boundaries given by the conditions $\Delta M^{\prime}(Q_{\text{max}})=0$ and $R^{\prime}(Q_{\text{min}})=0$. Using Eq. (31), find that $\lambda_{\text{max}}=\infty$ and $\lambda_{\text{min}}=0$, and using Eq. (28), find that $\mu_{\text{max}}=\Delta M(Q_{\text{max}})$ and $\mu_{\text{min}}=\Delta M(Q_{\text{min}})$. Finally, summarize the solution for the different values of $\mu$. It is not possible for $\mu$ to be greater than $\mu_{\text{max}}$. If $\mu=\mu_{\text{max}}$
@@
The electronic currency management firm will want to maximize $\mu$ in order to maximize the value of the currency, so that $\mu=\mu_{\text{max}}$. Therefore, the solution is given by Eq. (32). For a well managed currency, $m_{e}S_{0}/T_{I}\gg 1$. For instance, the example given in App. A has $T_{I}=1\,\text{yr}$, $S_{0}=2\,\text{yr}$ and $m_{e}=3.5$, so that $m_{e}S_{0}/T_{I}=7$ and $R^{\prime}=(1/8)E^{\prime}$. This is close to being a revenue maximizing firm with $R^{\prime}=0$ and $\mu=\mu_{\text{min}}$, which is not much less than $\mu_{\text{max}}$ since
@@
for the example given in App. A. In contrast, if the local approximation is made, the value of the currency is still maximized by using Eq. (32), but the non-local business network and time effects are neglected. The effective economic multiplier $m_{e}$ will approach 1, and the savings multiplier $S_{0}$ will approach 0. For Target Energy Inc., the typical firm operated with the local assumption discussed in App. A, $m_{e}=2$ and $S_{0}=1/20\,\text{yr}$. This gives $m_{e}S_{0}/T_{I}=1/10$, so that $R^{\prime}=(10/11)E^{\prime}\approx E^{\prime}$ – a profit, not revenue, maximizing firm.
The functional integration over the distribution given by the functional $f[R(t)]$ in Eq. (9) is how the risk model enters this theory. This is what could be approximated by the Generative Pretrained Transformer (GPT) (Radford et al., 2018; Vaswani, 2017; Li et al., 2022), a Generative Adversarial Network (GAN) (Goodfellow et al., 2016), the Mallat Scattering Transformation (MST) (Glinsky and Maupin, 2023; Glinsky, 2024a), or the Heisenberg Scattering Transformation (HST) (Glinsky, 2024b, c, 2023, d, e), all forms of Artificial Intelligence (AI) (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016). In the case of the local approximation, it is approximated by solving the diffusive Black-Scholes equation given in Eq. (6). There are also more simple approximations that can be made that respect the low risk and long-term nature of many financial situations.
Starting with the equation for $R_{0}$, Eq. (9), and substituting the solutions to the Black-Scholes equation, Eq. (6), into it, one can derive the expression for NPV given in Eq. (1). It will be shown in Sec. VII and Sec. IX how a constrained optimization of the expression for $R_{0}$ given in Eq. (9) leads to the optimization of NPV given in Eq. (1), when a dominate diffusive risk model is superimposed on the conservative financial dynamics of a sub-economy.
The dramatic difference between using the true distribution functional $f[R(t)]$ of Eq. (9) and the distribution function $f(J,t)$ of Eq. (6) is demonstrated in Fig. 3. In this figure the HST and the methods of AI are used to conditionally generate the distribution based on the historical behavior of the oil price (as will be explained in detail in Sec. VII). Note the discounting of the value of the next business cycle by the local approximation, so that no inventory would be amassed to take advantage of the next market upswing. Also note how poor the simulations using the local approximations are – that is how bad the approximation is.
Figure 3: Examples of conditional generation of price profile $R(t)$. (a) using the distribution functional $f[R(t)]$ of Eq. (9) via the HST, (b) using the distribution function $f(R,t)$ of Eq. (6). The thick black line is the actual signal. The thin black lines are the conditional realizations.
@@
In Sec. VII, it will be shown how to solve Eq. (37) using the HST and the methods of AI based on either simulations of the system or observations of the system, driven by an external force $F_{\text{ext}}$. The solution will be the fiscal (investment and operational) and monetary policy $\pi^{*}(q)$ that will locally maximize $V(q,t)$, but needs to be stabilized and cooled since local maximums are unstable equilibriums. The reward that is optimized $R(q)$ can be modified to take into account other benefits or costs to society, such as sustainability and beauty, by applying a conservative economic control force $\Delta p/\Delta t=F_{c}(q)$. Furthermore, the application of the economic force (arbitrage trading) to stabilize the system equilibrium $\pi^{*}(q)$ and minimize the fluctuations about the system equilibrium can be done in a similarly direct feedback manner by applying a force $F_{\text{sf}}(P)$ given in Eq. (99). It also can be done by applying a ponderomotive and diffusive force $F_{\text{sp}}(t)$, given in Eq. (100), that is not dependant on knowing the state of the system. More details of these control forces can be found in Sec. VII.
If instead of the true model of the system, the diffusive model of Eq. (6) is used and applied to solve Eq. (37), it can be shown that the equilibrium optimum policy $\pi^{*}(q)$ is to optimize the NPV of Eq. (1) and to invest with terms of debt repayment. This is equivalent to the profit maximization of the previous constrained optimization, instead of a revenue maximization. It should be noted that the local NPV maximization leads to a significant reduction in the long-term economic activity and even long-term profits. This greedy, local optimization when the system does not have exponential risk (that is diffusive dynamics), leads to very poor strategy (that is, policy). It would be like playing chess with no regard for the long-term effects of the next move. One would never sacrifice a piece to improve the long-term prospects of winning. One would only consider what pieces could be captured or lost with the next move or two. It also would be like playing PacMan with a strategy to swallow the next cherry as quickly as possible with no concern to whether that path leads to being consumed by a ghost in the near future.
@@ V Management of sub-economy
The criteria on which investments should be made is whether the demand for the electronic currency that the investment will create is greater than the investment, as shown in Eq. (22). The electronic currency management firm can then electronic currency leverage the funds that it has in reserves by issuing new electronic currency to make the investment. This ensures a future growth in the value of the currency, and that the sub-economy will not be paying for the investment through inflation, effectively an inflation tax. This is in contrast to issuing debt, which exploits the sub-economy and leads to sub-optimal levels of investment and economic performance of the sub-economy.
@@ VI Sub-economy as a nonlinear system
In this section, we will approach the understanding of a financial system as a nonlinear system like Complexity Economics (Farmer, 2024) does, and the financial and monetary policies as a problem in nonlinear systems control. We start this discussion by referring to Fig. 5, which is a representative contour plot for a physical system (Kuzmin et al., 2004). It also can be looked at as a topography map. This map has two basins with basin centers indicated by the o-points, and one saddle point (mountain pass) between the basins indicated by the x-point. There are strong exploitative thermal forces that will take a sub-economy to the o-point once in the respective basin. They are also what are called stable equilibriums. The discipline of nonlinear dynamics refers to these o-points or stable equilibriums as attractors. This is in contrast to the x-point, which is the point that the sub-economy descends to from the mountain, but it is not a stable equilibrium so that as it is approached the sub-economy will fall into one of the exploitative basins. The discipline of nonlinear dynamics refers to these x-points or unstable/metastable equilibriums as semi-attractors (that is half attractor and half repeller). Semi-attractors first attract the trajectory to them, but once reached repel the trajectory away from them. These unstable equilibriums are the desirable states from a social aesthetic perspective as shown in Fig. 6. They need an active control system, though, in order to stabilize them. This is like stabilizing an inverted pendulum with a vibrating saw as shown in Fig. 7 (Kapitza, 1951a, b; Landau and Lifshitz, 1976). Stabilization is the construction of a small alpine valley at the mountain pass where it is easiest to do. The electronic currency management firm needs to vibrate the market with its arbitrage trading to stabilize the system. The technical problem is in identifying the natural frequencies of the system. For instance, when one flies a plane, one must take into account that it takes several seconds for the plane to respond. If one tries to make corrections faster than this, one will over-correct and cause the plane to go out of control. The issue with financial systems is that they have many natural frequencies, in fact an infinite number. The electronic currency management firm needs to use innovative genAI technology that will control the financial system at all the natural frequencies. Details of how this is done are given in the US Patent Application, “Systems and Methods for Controlling Complex Systems” (Glinsky, 2024f) and discussed in Sec. VII. Current control systems only have a single time scale, and often have catastrophic phase lags built in that destabilize the system like proof-of-stake systems (e.g., New Ethereum) and the buying and selling of bonds by a central bank (e.g., US T-bills and British GILTs).
@@
Figure 7: Stabilization of the pendulum about the unstable equilibrium (saddle point) by use of nonlinear control. The full YouTube video can be found at this Link.
Note that the electronic currency management firm will want to buy the electronic currency, using its cash reserves, when the price is less than the equilibrium price to increase the demand and therefore the price, and sell the electronic currency when the price is more than the equilibrium price to increase the supply and therefore decrease the price. That is, it will buy low and sell high – a money making financial heat engine. Mother nature rewards doing what she wants. The essential tricks are sensing what the equilibrium is and doing it on multiple time scales. It is all about topological discovery of the financial system response and equilibrium, then topological manipulation to control and stabilize the financial system (Glinsky, 2024d, e). This is not the case today.